# SimROD: A Simple Baseline for Raw Object Detection with Global and Local Enhancements

### v2

@@ Abstract
Most visual models are designed for sRGB images, yet RAW data offers significant advantages for object detection by preserving sensor information before ISP processing. This enables improved detection accuracy and more efficient hardware designs by bypassing the ISP. However, RAW object detection is challenging due to limited training data, unbalanced pixel distributions, and sensor noise. To address this, we propose SimROD, a lightweight and effective approach for RAW object detection. We introduce a Global Gamma Enhancement (GGE) module, which applies a learnable global gamma transformation with only four parameters, improving feature representation while keeping the model efficient. Additionally, we leverage the green channel’s richer signal to enhance local details, aligning with the human eye’s sensitivity and Bayer filter design. Extensive experiments on multiple RAW object detection datasets and detectors demonstrate that SimROD outperforms state-of-the-art methods like RAW-Adapter and DIAP while maintaining efficiency. Our work highlights the potential of RAW data for real-world object detection. Our code is available on our project page: https://ocean146.github.io/SimROD2025/.
1 Introduction
Accurate object detection is crucial for autonomous driving, especially under challenging lighting and weather conditions. Traditional methods that rely on sRGB images often lose important details during processing. In contrast, RAW sensor data captures the unprocessed, richer signal from the sensor, preserving more details and a wider dynamic range [39, 8, 23, 7]. Moreover, as shown in Figure  1(a), by directly using RAW data, there’s no need for an ISP module, which can reduce system complexity, lower latency, and cut costs—key benefits for lightweight, real-time applications.
@@ 1 Introduction
However, working with RAW data introduces several challenges, including limited training samples, unbalanced pixel distributions, and sensor noise. Current approaches typically rely on complex frameworks that integrate end-to-end optimization of ISP stages with object detection models. These methods explicitly designs learnable ISP stages to transform RAW data [39, 23, 24, 42]. While these methods demonstrate promising results, they tend to be computationally expensive and introduce unnecessary design complexities. Additionally, modern cameras emphasize the green channel in their Bayer filter design [47, 1] as the human eye is highly sensitive to green light in both bright and low-light conditions [33]. However, most existing methods treat RGB channels equally, overlooking the green channel’s unique advantages in RAW data.
In this work, we present SimROD, a simple yet effective approach that enhances RAW object detection performance while maintaining model simplicity. Our approach is based on two key insights: (1) learning an adapted global transformation might not be complicated but crucial for fine-grained tasks [39, 4], and (2) the superior informativeness of the green channel (Figure 1(b)) in the RGGB Bayer pattern. By leveraging these insights, we introduce an efficient Global Gamma Enhancement (GGE) with only four learnable parameters, significantly reducing model complexity while achieving comparable performance to more complex methods. We also propose a Green-Guided Local Enhancement (GGLE) module that uses the green channel to refine local image details, further boosting detection accuracy.
Through extensive experiments, we demonstrate that SimROD outperforms existing methods, such as RAW-Adapter [8] and DIAP [39], on several standard RAW object detection benchmarks, including ROD [39], LOD [15], and Pascal-Raw [27]. For example, on the benchmark of Pascal-Raw [27], following the setup of RAW-Adapter [8], we achieve consistent performance improvement across different object detectors and different setups of Pascal-Raw [27]. Furthermore, we create a strong baseline for DIAP [39] by leveraging weights pre-trained on MS COCO [20], which raises its performance from 24.0% mAP to 30.7% mAP on the ROD dataset [39]. SimROD achieves notable improvements even relative to this strong baseline.
@@
  • Despite its simplicity, SimROD achieves state-of-the-art performance on ROD [39], LOD [15], and Pascal-Raw [27], surpassing prior methods like RAW-Adapter [8] and DIAP [39].
Our code is available on: https://ocean146.github.io/SimROD2025/.
2 Motivation
The human visual system exhibits a pronounced sensitivity to green light wavelengths in both bright and dim light conditions, as evidenced by [33]. Due to this characteristic of human vision, cameras prioritize green channels in their Bayer filter design [1, 47]. Motivated by this biological and technical precedent, we explored the effectiveness of green channel for object detection by analyzing channel sensitivity and signal-to-noise ratio (SNR) for individual channels.
Figure 2: Left: We evaluate RAW object detection on the LOD dataset [15] using individual color channels—green (G), red (R), and blue (B)—with the state-of-the-art DIAP method [39]. The results highlight the superior performance of G. Right: G has a significantly higher SNR than R and B, suggesting it may be more resistant to noise in extreme lighting conditions, potentially improving robustness.
  • Channel Sensitivity Analysis. We utilized DIAP [39] on the LOD [15] dataset to independently assess the detection performance of each channel (green, red, and blue). As depicted in Figure 2 left, the green channel achieved the highest detection accuracy, surpassing the red and blue channels by substantial margins (approximately 10 and 20 AP points, respectively), underscoring its superior informativeness for object detection in RAW data.
@@ 3 Method
The proposed method’s overall framework is illustrated in Figure 3. A RAW image is an unprocessed digital photograph that retains all the data from a camera’s sensor, including the RGGB color pattern. Given a RAW image $X_{\text{RAW}}\in\mathbb{R}^{2H\times 2W}$, we repack and convert it into a four-channel image $X_{\text{packed}}\in\mathbb{R}^{H\times W\times 4}$, where the last dimension represents the color channels in the RGGB pattern.
@@
This section is organized as follows: in Section 3.1, we provide a detailed description of the proposed the Global Gamma Enhancement (GGE) module; in Section 3.2, we detail the module Green-Guided Local Enhancement (GGLE); implementation details are provided in Section 3.3.
3.1 Global Gamma Enhancement
In visual perception tasks involving raw sensor data, pixel values are typically concentrated in the low range, which makes it challenging for deep neural networks to learn effectively and extract useful features [39]. Therefore, dynamic range adjustment is an essential step in image signal processing pipelines to prepare raw data for object detection [39, 4].
@@ Discussion
In contrast to the recent method introduced by [39], which predicts parameters for image-level RAW data adjustment, our proposed GGE contains only four parameters, thus the network only consists of a minimal number of parameters. This results in greater computational efficiency while achieving comparable or superior performance to [39] (Section 4.4). Notably, we observed that the gamma parameters predicted by the image-level adjustment [39] module remained largely unchanged, even when a completely random noise image was used as input.
3.2 Green-Guided Local Enhancement
@@
As illustrated in Figure 3, GGLE consists of two primary branches. The first branch, the RGGB branch, processes the complete RGGB data $X_{\gamma}$ using a convolutional neural network, $\mathcal{F}_{l}$, which extracts spatial features from all channels to produce a feature map, $\mathcal{F}_{l}(X_{\gamma})$, representing the full image context. The second branch, the Green branch, specifically targets the two green channels, $X_{\gamma}^{G_{1}}$ and $X_{\gamma}^{G_{2}}$, which are concatenated and processed through another convolutional network, $\mathcal{F}_{l}^{G}$, resulting in a green-focused feature map, $\mathcal{F}_{l}^{G}(X_{\gamma}^{G})$, where $X_{\gamma}^{G}=[X_{\gamma}^{G_{1}},X_{\gamma}^{G_{2}}]$.
@@
3.3 Implementation Details
In the Global Gamma Enhancement (GGE) module, each $\gamma_{i}$ is parameterized in a straightforward manner. For each $\gamma_{i}$, we define a learnable parameter $\alpha_{i}\in\mathbb{R}$, which is constrained to the range $(-1,1)$ using the tanh activation function. This output is then linearly scaled to lie within a predefined range $(\gamma_{\text{min}},\gamma_{\text{max}})$, where $\gamma_{\text{min}}$ and $\gamma_{\text{max}}$ are hyperparameters. Following the settings in [39], we set $\gamma_{\text{max}}=1/7.0$ and $\gamma_{\text{min}}=1/10.5$ by default. For the over-exposure datasets ADE20k-Raw [8] and PascalRAW [8], $\gamma_{\text{max}}$ is set to $1/0.5$ to facilitate the restoration of over-exposed images.
For the Green-Guided Local Enhancement (GGLE) module, both $\mathcal{F}_{l}$ and $\mathcal{F}_{l}^{G}$, used in the RGGB and Green branches respectively, employ a simple yet effective architecture consisting of convolutional layers, Batch Normalization [16], and LeakyReLU [25] activation functions. Collectively, GGE and GGLE comprise a total of just 0.003 million parameters, rendering these modules lightweight compared to previous approaches. In contrast, prior methods often require hundreds of times more parameters, yet our approach achieves superior performance.
@@ 3.3 Implementation Details
Our SimROD is an end-to-end framework, jointly optimizing the GGE and GGLE modules alongside the downstream model, thereby eliminating the need for additional loss functions tailored to these enhancement stages. Furthermore, we adopt the same loss functions as those employed in the original works [44, 21, 37]. For example, when using YoloX [44] as the detector, optimization relies solely on the standard detection loss, which includes classification and regression components. The total loss function is defined as:
@@
where $\lambda$ is the default balancing factor of the detector between the classification loss $L_{cls}$ and the regression loss $L_{reg}$ ($\lambda=3$ in YoloX [44]). This unified approach allows the enhancement modules to adapt naturally to the detection objectives, supporting end-to-end optimization of the entire framework.
4 Experiments
@@
Method LOD [15] Pascal-Raw [27] Add.
AP AP50 AP AP50 Params (M)
Demosacing [8] 58.5 58.5 89.2 89.2 0.000
Default ISP [27, 8] 58.4 58.4 89.6 89.6 0.000
Karaimer et Brown [18, 8] 54.4 54.4 89.4 89.4 -
Lite-ISP [43, 8] - - 88.5 88.5 -
InvISP [38, 8] 56.9 56.9 87.6 87.6 -
Dirty-Pixel [9, 8] 61.6 61.6 89.7 89.7 4.28
[…]
RAW-Adapter [8] 62.1 62.1 89.7 89.7 0.46
4.1 Datasets and Evaluation Metrics
@@
We evaluate and compare our method against existing methods on four benchmark datasets: Pascal-Raw [27], LOD [15], and ROD [39] for object detection, and ADE20K-Raw [8] for semantic segmentation.
Pascal-Raw [27].
@@
The Pascal-Raw dataset [27] comprises 4,259 RAW images captured under standard lighting conditions with a Nikon D3200 DSLR camera, covering three object classes: person, car, and bicycle. The dataset is divided into 2,129 images for training and 2,130 images for testing, as defined by [27]. For our experiments, we utilize the preprocessed RAW data provided by RAW-Adapter [8]. Additionally, we evaluate the Over-exp and Dark versions of Pascal-Raw [27], generated by RAW-Adapter [8] using InvISP [38], which include challenging overexposed and dark scenes, respectively.
LOD [15].
@@
The LOD dataset [15] contains 2,230 RAW images captured with a Canon EOS 5D Mark IV camera in low-light conditions, covering eight object classes: bus, chair, TV monitor, bicycle, bottle, dining table, motorbike, and car. The dataset is split into 1,800 images for training and 430 images for testing. For all our experiments, we use the preprocessed RAW data provided by RAW-Adapter [8] to ensure a fair comparison with RAW-Adapter.
ROD [39].
@@
The original ROD dataset [39] contains 25,207 RAW images, including 10k daytime scenes and 14k nighttime scenes, across six common object categories. Compared to other datasets, ROD is a larger-scale dataset with a more diverse range of scenes, focusing on urban driving scenarios. Due to limitations in dataset access, we were unable to adhere to the “standard” partitioning protocol defined by ROD [39]. Despite repeated attempts to obtain the full dataset, only a subset of the training data has been publicly released. This subset consists of 16,089 RAW images, including 4,053 daytime scenes and 12,036 nighttime scenes, but it contains only five object classes instead of the six originally specified in [39]. To ensure a fair evaluation, we randomly partitioned the publicly available subset into 80% for training and 20% for testing. This partition resulted in 3,245 daytime scenes and 9,626 nighttime scenes in the training set, with the remaining images reserved for testing. All references to the ROD dataset in this paper pertain specifically to this re-partitioned subset. We will make this subset publicly available to facilitate the re-implementation of our results.
ADE20K-Raw [8].
@@
The ADE20K-RAW dataset [8] is a RAW-format segmentation dataset derived from ADE20K [45], consisting of 27,574 images synthesized using InvISP [38] by RAW-Adapter [8]. This dataset includes three versions—normal, dark, and over-exposure—that simulate different lighting conditions. The dataset follows the same training and testing splits defined by ADE20K [45] for consistency.
@@
4.2 Training Details
@@
For detection tasks, we conduct experiments using two object detectors: YoloX [44] and RetinaNet [21], following the protocols of DIAP [39] and RAW-Adapter [8].
@@
For YoloX [44], we use the official training strategies, including standard data augmentation techniques such as random horizontal flipping, scale jittering through resizing, and Mosaic augmentation [2]. Both training and testing data are resized to 640×640. The model is trained for 300 epochs, with five warmup epochs, using the SGD optimizer with a momentum of 0.9. We apply a cosine learning rate schedule and use a batch size of 12. The training process of Pascal-Raw [27] and LOD [15] takes 2 hours on three NVIDIA RTX 3090 GPUs. The training process of ROD [39] takes 10 hours on three NVIDIA RTX 4090 GPUs. For initialization, we use COCO [20] pre-trained weights, which improve the performance from 24.0% AP to 30.7% AP compared to the approach proposed in the original DIAP [39] and construct a strong baseline.
@@
For the RetinaNet [21] detector, we use the MMDetection framework and its default data augmentation pipeline, which includes random cropping, random flipping, and multi-scale testing [8]. We set the learning rate for the proposed SimROD to 3e-3.
@@
For the segmentation task, we followed the settings from RAW-Adapter [8] and used Segformer [37] with the MIT-B5 [37] backbone. We trained the model on four NVIDIA RTX 4090 GPUs, with learning rates adjusted empirically for the proposed method: 8e-5 for the normal-light dataset of Pascal-Raw [8], 7e-5 for the dark dataset of Pascal-Raw [8], and 9e-5 for the over-exposure dataset [8].
4.3 Comparison with State-of-the-Art Methods
@@
RAW Object Detection on Pascal-Raw [27], LOD [15], and ROD [39] Datasets.
@@
Table 1 and Table 2 demonstrate that our proposed SimROD consistently outperforms existing methods across all datasets while being highly efficient. To ensure a fair comparison with RAW-Adapter [8] and DIAP [39], we strictly follow their official settings with YoloX-Tiny [44] and RetinaNet-R50 [21]. For YoloX-Tiny (Table 1), SimROD achieves notable AP gains of +2.4 on ROD, +0.8 on LOD, and +1.0 on Pascal-Raw, surpassing DIAP [39] with just 0.003M additional parameters. For RetinaNet-R50 (Table 2), SimROD reaches an impressive 63.4% AP50 on LOD, outperforming RAW-Adapter [8] and DIAP [39], while requiring only 0.7% of RAW-Adapter’s additional parameters. These results highlight SimROD’s strong generalization, superior accuracy, and exceptional parameter efficiency, making it a highly effective solution for RAW object detection.
Semantic Segmentation on ADE20K-Raw [45, 8] with RAW Data.
@@
To further validate SimROD’s effectiveness, we evaluate it on semantic segmentation using ADE20K-Raw with Segformer [37], following RAW-Adapter [8]. As shown in Table 3, SimROD achieves the best performance across normal and low-light conditions, while achieving competitive performance on over-exposure conditions, demonstrating strong generalization. Despite adding only 0.003M parameters, it remains highly efficient while delivering superior accuracy. These results reinforce our object detection findings, proving SimROD’s versatility in RAW data processing. Its potential extends beyond detection, making it applicable to broader vision tasks.

### v3

@@ Abstract
Most visual models are designed for sRGB images, yet RAW data offers significant advantages for object detection by preserving sensor information before ISP processing. This enables improved detection accuracy and more efficient hardware designs by bypassing the ISP. However, RAW object detection is challenging due to limited training data, unbalanced pixel distributions, and sensor noise. To address this, we propose SimROD, a lightweight and effective approach for RAW object detection. We introduce a Global Gamma Enhancement (GGE) module, which applies a learnable global gamma transformation with only four parameters, improving feature representation while keeping the model efficient. Additionally, we leverage the green channel’s richer signal to enhance local details, aligning with the human eye’s sensitivity and Bayer filter design. Extensive experiments on multiple RAW object detection datasets and detectors demonstrate that SimROD outperforms state-of-the-art methods like RAW-Adapter and DIAP while maintaining efficiency. Our work highlights the potential of RAW data for real-world object detection.
Project pagehttps://ocean146.github.io/SimROD2025/
Extended versionhttps://arxiv.org/abs/2503.07101
Introduction
Accurate object detection is crucial for autonomous driving, especially under challenging lighting and weather conditions. Traditional methods that rely on sRGB images often lose important details during processing. In contrast, RAW sensor data captures the unprocessed, richer signal from the sensor, preserving more details and a wider dynamic range (Xu et al. 2023; Cui and Harada 2024; Morawski et al. 2022; Chen, Tai, and Ma 2024). Moreover, as shown in Figure  1(a), by directly using RAW data, there’s no need for an ISP module, which can reduce system complexity, lower latency, and cut costs—key benefits for lightweight, real-time applications.
@@ Introduction
However, working with RAW data introduces several challenges, including limited training samples, unbalanced pixel distributions, and sensor noise. Current approaches typically rely on complex frameworks that integrate end-to-end optimization of ISP stages with object detection models. These methods explicitly designs learnable ISP stages to transform RAW data (Xu et al. 2023; Morawski et al. 2022; Mosleh et al. 2020; Yu et al. 2021). While these methods demonstrate promising results, they tend to be computationally expensive and introduce unnecessary design complexities. Additionally, modern cameras emphasize the green channel in their Bayer filter design (Zou, Yan, and Fu 2023; Bayer 1976) as the human eye is highly sensitive to green light in both bright and low-light conditions (Wald 1945). However, most existing methods treat RGB channels equally, overlooking the green channel’s unique advantages in RAW data.
In this work, we present SimROD, a simple yet effective approach that enhances RAW object detection performance while maintaining model simplicity. Our approach is based on two key insights: (1) learning an adapted global transformation might not be complicated but crucial for fine-grained tasks (Xu et al. 2023; Buckler, Jayasuriya, and Sampson 2017), and (2) the superior informativeness of the green channel (Figure 1(b)) in the RGGB Bayer pattern. By leveraging these insights, we introduce an efficient Global Gamma Enhancement (GGE) with only four learnable parameters, significantly reducing model complexity while achieving comparable performance to more complex methods. We also propose a Green-Guided Local Enhancement (GGLE) module that uses the green channel to refine local image details, further boosting detection accuracy.
Through extensive experiments, we demonstrate that SimROD outperforms existing methods, such as RAW-Adapter (Cui and Harada 2024) and DIAP (Xu et al. 2023), on several standard RAW object detection benchmarks, including ROD (Xu et al. 2023), LOD (Hong et al. 2021), and Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014). For example, on the benchmark of Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014), following the setup of RAW-Adapter (Cui and Harada 2024), we achieve consistent performance improvement across different object detectors and different setups of Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014). Furthermore, we create a strong baseline for DIAP (Xu et al. 2023) by leveraging weights pre-trained on MS COCO (Lin et al. 2014), which raises its performance from 24.0% mAP to 30.7% mAP on the ROD dataset (Xu et al. 2023). SimROD achieves notable improvements even relative to this strong baseline.
@@
  • Despite its simplicity, SimROD achieves state-of-the-art performance on ROD (Xu et al. 2023), LOD (Hong et al. 2021), and Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014), surpassing prior methods like RAW-Adapter (Cui and Harada 2024) and DIAP (Xu et al. 2023).
Motivation
The human visual system exhibits a pronounced sensitivity to green light wavelengths in both bright and dim light conditions, as evidenced by (Wald 1945). Due to this characteristic of human vision, cameras prioritize green channels in their Bayer filter design (Bayer 1976; Zou, Yan, and Fu 2023). Motivated by this biological and technical precedent, we explored the effectiveness of green channel for object detection by analyzing channel sensitivity and signal-to-noise ratio (SNR) for individual channels.
Figure 2: Left: We evaluate RAW object detection on the LOD dataset (Hong et al. 2021) using individual color channels—green (G), red (R), and blue (B)—with the state-of-the-art DIAP method (Xu et al. 2023). The results highlight the superior performance of G. Right: G has a significantly higher SNR than R and B, suggesting it may be more resistant to noise in extreme lighting conditions, potentially improving robustness.
  • Channel Sensitivity Analysis. We utilized DIAP (Xu et al. 2023) on the LOD (Hong et al. 2021) dataset to independently assess the detection performance of each channel (green, red, and blue). As depicted in Figure 2 left, the green channel achieved the highest detection accuracy, surpassing the red and blue channels by substantial margins (approximately 10 and 20 AP points, respectively), underscoring its superior informativeness for object detection in RAW data.
@@
The proposed method’s overall framework is illustrated in Figure 3. A RAW image is an unprocessed digital photograph that retains all the data from a camera’s sensor, including the RGGB color pattern. Given a RAW image $(X_{\text{RAW}}\in\mathbb{R}^{2H\times 2W}$, we repack and convert it into a four-channel image $X_{\text{packed}}\in\mathbb{R}^{H\times W\times 4}$, where the last dimension represents the color channels in the RGGB pattern.
@@
This section is organized as follows: in Section Global Gamma Enhancement, we provide a detailed description of the proposed the Global Gamma Enhancement (GGE) module; in Section Green-Guided Local Enhancement, we detail the module Green-Guided Local Enhancement (GGLE); implementation details are provided in Section Implementation Details.
Global Gamma Enhancement
In visual perception tasks involving raw sensor data, pixel values are typically concentrated in the low range, which makes it challenging for deep neural networks to learn effectively and extract useful features (Xu et al. 2023). Therefore, dynamic range adjustment is an essential step in image signal processing pipelines to prepare raw data for object detection (Xu et al. 2023; Buckler, Jayasuriya, and Sampson 2017).
@@ Discussion
In contrast to the recent method introduced by (Xu et al. 2023), which predicts parameters for image-level RAW data adjustment, our proposed GGE contains only four parameters, thus the network only consists of a minimal number of parameters. This results in greater computational efficiency while achieving comparable or superior performance to (Xu et al. 2023) (Section Analysis and Discussion). Notably, we observed that the gamma parameters predicted by the image-level adjustment (Xu et al. 2023) module remained largely unchanged, even when a completely random noise image was used as input.
Green-Guided Local Enhancement
@@
As illustrated in Figure 3, GGLE consists of two primary branches. The first branch, the RGGB branch, processes the complete RGGB data $X_{\gamma}$ using a convolutional neural network, $\mathcal{F}_{l}$, which extracts spatial features from all channels to produce a feature map, $\mathcal{F}_{l}(X_{\gamma})$, representing the full image context. The second branch, the Guidance branch, specifically targets the two green channels, $X_{\gamma}^{G_{1}}$ and $X_{\gamma}^{G_{2}}$, which are concatenated and processed through another convolutional network, $\mathcal{F}_{l}^{G}$, resulting in a green-focused feature map, $\mathcal{F}_{l}^{G}(X_{\gamma}^{G})$, where $X_{\gamma}^{G}=[X_{\gamma}^{G_{1}},X_{\gamma}^{G_{2}}]$.
@@
Implementation Details
In the Global Gamma Enhancement (GGE) module, each $\gamma_{i}$ is parameterized in a straightforward manner. For each $\gamma_{i}$, we define a learnable parameter $\alpha_{i}\in\mathbb{R}$, which is constrained to the range (1,1)(-1,1) using the tanh activation function. This output is then linearly scaled to lie within a predefined range $(\gamma_{\text{min}},\gamma_{\text{max}})$, where $\gamma_{\text{min}}$ and $\gamma_{\text{max}}$ are hyperparameters. Following the settings in (Xu et al. 2023), we set $\gamma_{\text{max}}=1/7.0$ and $\gamma_{\text{min}}=1/10.5$.
For the Green-Guided Local Enhancement (GGLE) module, both $\mathcal{F}_{l}$ and $\mathcal{F}_{l}^{G}$, used in the RGGB and Guidance branches respectively, employ a simple yet effective architecture consisting of convolutional layers, Batch Normalization (Ioffe and Szegedy 2015), and LeakyReLU (Nair and Hinton 2010) activation functions. Collectively, GGE and GGLE comprise a total of just 0.003 million parameters, rendering these modules lightweight compared to previous approaches. In contrast, prior methods often require hundreds of times more parameters, yet our approach achieves superior performance.
@@
Our SimROD is an end-to-end framework, jointly optimizing the GGE and GGLE modules alongside the downstream model, thereby eliminating the need for additional loss functions tailored to these enhancement stages. Furthermore, we adopt the same loss functions as those employed in the original works (Zheng et al. 2021; Lin et al. 2017; Xie et al. 2021). For example, when using YoloX (Zheng et al. 2021) as the detector, optimization relies solely on the standard detection loss, which includes classification and regression components. The total loss function is defined as:
@@
where $\lambda$ is the default balancing factor of the detector between the classification loss $L_{cls}$ and the regression loss $L_{reg}$ ($\lambda=3$ in YoloX (Zheng et al. 2021)). This unified approach allows the enhancement modules to adapt naturally to the detection objectives, supporting end-to-end optimization of the entire framework.
Experiments
Table 1: Results with YoloX-Tiny (Zheng et al. 2021), following DIAP’s benchmark (Xu et al. 2023). Performance metrics (AP and AP50) for YoloX-Tiny across different methods including IA (Xu et al. 2023), Raw-or-cook (Ljungbergh et al. 2023), GenISP (Morawski et al. 2022), RAW-Adapter (Cui and Harada 2024) and DIAP (Xu et al. 2023). The best performance for each dataset is highlighted in bold. The table also includes the number of parameters (in millions). The Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014) in this table is the normal light version. $\dagger$} indicates reproduced results. N/A means the model did not converge.
Method LOD Pascal-Raw ROD Add.
AP AP50 AP AP50 AP AP50 Params (M)
Demosacing 26.5 46.0 66.8 92.8 4.8 9.5 0.000
Gamma 25.7 44.2 69.0 94.5 7.6 14.3 0.000
IA 25.1 43.9 68.9 94.2 30.6 53.3 0.176
@@
DIAP 25.9 43.4 68.7 94.2 30.7 53.4 0.260
Our SimROD 26.7(+0.8) 46.3(+2.9) 69.7(+1.0) 95.1(+1.1) 33.1(+2.4) 57.6(+4.2) 0.003 (1%)
Datasets and Evaluation Metrics
@@
We evaluate and compare our method against existing methods on four benchmark datasets: Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014), LOD (Hong et al. 2021), and ROD (Xu et al. 2023) for object detection, and ADE20K-Raw (Cui and Harada 2024) for semantic segmentation.
Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014).
@@
The Pascal-Raw dataset (Omid-Zohoor, Ta, and Murmann 2014) comprises 4,259 RAW images captured under standard lighting conditions with a Nikon D3200 DSLR camera, covering three object classes: person, car, and bicycle. The dataset is divided into 2,129 images for training and 2,130 images for testing. For our experiments, we utilize the preprocessed RAW data provided by RAW-Adapter (Cui and Harada 2024).
LOD (Hong et al. 2021).
@@
The LOD dataset (Hong et al. 2021) contains 2,230 RAW images captured with a Canon EOS 5D Mark IV camera in low-light conditions, covering eight object classes: bus, chair, TV monitor, bicycle, bottle, dining table, motorbike, and car. The dataset is split into 1,800 images for training and 430 images for testing. For all our experiments, we use the preprocessed RAW data provided by RAW-Adapter (Cui and Harada 2024) to ensure a fair comparison with RAW-Adapter.
ROD (Xu et al. 2023).
@@
The original ROD dataset (Xu et al. 2023) contains 25,207 RAW images, including 10k daytime scenes and 14k nighttime scenes, across six common object categories. Compared to other datasets, ROD is a larger-scale dataset with a more diverse range of scenes, focusing on urban driving scenarios. Due to limitations in dataset access, we were unable to adhere to the “standard” partitioning protocol defined by ROD (Xu et al. 2023). Despite repeated attempts to obtain the full dataset, only a subset of the training data has been publicly released. This subset consists of 16,089 RAW images, including 4,053 daytime scenes and 12,036 nighttime scenes, but it contains only five object classes instead of the six originally specified in (Xu et al. 2023). To ensure a fair evaluation, we randomly partitioned the publicly available subset into 80% for training and 20% for testing. This partition resulted in 3,245 daytime scenes and 9,626 nighttime scenes in the training set, with the remaining images reserved for testing. We strictly followed the official guidelines for data preprocessing and conducted multiple experimental trials to ensure the validity of the results. All references to the ROD dataset in this paper pertain specifically to this re-partitioned subset. We will make this subset publicly available to facilitate the re-implementation of our results.
ADE20K-Raw (Cui and Harada 2024).
@@
The ADE20K-RAW dataset (Cui and Harada 2024) is a RAW-format segmentation dataset derived from ADE20K (Zhou et al. 2017), consisting of 27,574 images synthesized using InvISP (Xing, Qian, and Chen 2021) by RAW-Adapter (Cui and Harada 2024). This dataset includes three versions—normal, dark, and over-exposure—that simulate different lighting conditions. The dataset follows the same training and testing splits defined by ADE20K (Zhou et al. 2017) for consistency.
@@
Training Details
All experiments in this section follow the settings of DIAP (Xu et al. 2023), unless stated otherwise.
@@
For detection tasks, we conduct experiments using two object detectors: YoloX (Zheng et al. 2021) and RetinaNet (Lin et al. 2017), following the protocols of DIAP (Xu et al. 2023) and RAW-Adapter (Cui and Harada 2024). If not otherwise specified, all experiments were initialized with pre-trained weights.
@@
For YoloX (Zheng et al. 2021), we use the official training strategies, including standard data augmentation techniques such as random horizontal flipping, scale jittering through resizing, and Mosaic augmentation (Bochkovskiy, Wang, and Liao 2020). Both training and testing data are resized to 640×640. The model is trained for 300 epochs, with five warmup epochs, using the SGD optimizer with a momentum of 0.9. We apply a cosine learning rate schedule and use a batch size of 12. The training process of Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014) and LOD (Hong et al. 2021) takes 2 hours on three NVIDIA RTX 3090 GPUs. The training process of ROD (Xu et al. 2023) takes 10 hours on three NVIDIA RTX 4090 GPUs. For initialization, we use COCO (Lin et al. 2014) pre-trained weights, which improve the performance from 24.0% AP to 30.7% AP compared to the approach proposed in the original DIAP (Xu et al. 2023) and construct a strong baseline.
@@
For the RetinaNet (Lin et al. 2017) detector, we use the MMDetection framework and its default data augmentation pipeline, which includes random cropping, random flipping, and multi-scale testing (Cui and Harada 2024). We set the learning rate for the proposed SimROD to 3e-3.
@@
For the segmentation task, we followed the settings from RAW-Adapter (Cui and Harada 2024) and used Segformer (Xie et al. 2021) with the MIT-B5 (Xie et al. 2021) backbone. We trained the model on four NVIDIA RTX 4090 GPUs, with learning rates adjusted empirically for the proposed method: 8e-5 for the normal-light dataset of Pascal-Raw (Cui and Harada 2024), 7e-5 for the dark dataset of Pascal-Raw (Cui and Harada 2024), and 9e-5 for the over-exposure dataset (Cui and Harada 2024).
Comparison with State-of-the-Art Methods
Table 2: Results with RetinaNet-R50 (Lin et al. 2017), following RAW-Adapter’s benchmark (Cui and Harada 2024). Performance metrics (AP and AP50) for RetinaNet-R50 across different methods including Demosaicing (Cui and Harada 2024), Default ISP (Omid-Zohoor, Ta, and Murmann 2014), Karaimer et Brown (Karaimer and Brown 2016), Lite-ISP (Zhang et al. 2021), InvISP (Xing, Qian, and Chen 2021) and Dirty-Pixel (Diamond et al. 2021). The best performance for each dataset is highlighted in bold. The table also includes the number of parameters (in millions). The Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014) in this table is the normal light version. $\dagger$} indicates results reproduced using the official code. RetinaNet-R50 shows similar AP/AP50 due to strong performance and low small-object ratios in LOD/Pascal-Raw (5%/3% vs. 42% in COCO).
Method LOD Pascal-Raw Add.
AP AP50 AP AP50 Params (M)
Demosacing 58.5 58.5 89.2 89.2 0.000
Default ISP 58.4 58.4 89.6 89.6 0.000
Karaimer et Brown 54.4 54.4 89.4 89.4 -
Lite-ISP - - 88.5 88.5 -
InvISP 56.9 56.9 87.6 87.6 -
Dirty-Pixel 61.6 61.6 89.7 89.7 4.28
[…]
RAW-Adapter 62.1 62.1 89.7 89.7 0.46
@@
RAW Object Detection on Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014), LOD (Hong et al. 2021), and ROD (Xu et al. 2023) Datasets.
@@
Table 1 and Table 2 demonstrate that our proposed SimROD consistently outperforms existing methods across all datasets while being highly efficient. To ensure a fair comparison with RAW-Adapter (Cui and Harada 2024) and DIAP (Xu et al. 2023), we strictly follow their official settings with YoloX-Tiny (Zheng et al. 2021) and RetinaNet-R50 (Lin et al. 2017). For YoloX-Tiny (Table 1), SimROD achieves notable AP gains of +2.4 on ROD, +0.8 on LOD, and +1.0 on Pascal-Raw, surpassing DIAP (Xu et al. 2023) with just 0.003M additional parameters. For RetinaNet-R50 (Table 2), SimROD reaches an impressive 63.4% AP50 on LOD, outperforming RAW-Adapter (Cui and Harada 2024) and DIAP (Xu et al. 2023), while requiring only 0.7% of RAW-Adapter’s additional parameters. These results highlight SimROD’s strong generalization, superior accuracy, and exceptional parameter efficiency, making it a highly effective solution for RAW object detection.
Semantic Segmentation on ADE20K-Raw (Zhou et al. 2017; Cui and Harada 2024) with RAW Data.
@@
To further validate SimROD’s effectiveness, we evaluate it on semantic segmentation using ADE20K-Raw with Segformer (Xie et al. 2021), following RAW-Adapter (Cui and Harada 2024). As shown in Table 3, SimROD achieves the best performance across normal and low-light conditions, while achieving competitive performance on over-exposure conditions, demonstrating strong generalization. Despite adding only 0.003M parameters, it remains highly efficient while delivering superior accuracy. These results reinforce our object detection findings, proving SimROD’s versatility in RAW data processing. Its potential extends beyond detection, making it applicable to broader vision tasks.